OpenStudy (anonymous):

i need help please x/2 - (x-2/3)=0.4 solve for x

OpenStudy (sandra):

hey is the second part of the equation (x-2)/3, or is it x-(2/3) ?

OpenStudy (anonymous):

umm the equation is two fractions x/2 Subttact x-2/3equal 0.4

OpenStudy (sandra):

ok so x-2 is all on top of the second fraction?

OpenStudy (sandra):

\[\frac{x-2}{3}\]

OpenStudy (sandra):

like that ^ ?

OpenStudy (sandra):

so basically in that case you need to find a common denominator for x/2 and x-2/3

OpenStudy (sandra):

the common denominator would be 6 (the lowest number that 3 and 2 divide into evenly)

OpenStudy (anonymous):

umm the equation is two fractions x/2 Subttact x-2/3equal 0.4

OpenStudy (anonymous):

i got the answer now thanks

OpenStudy (sandra):

\[\frac{x}{2} - \frac{x-2}{3} = .4\] \[\frac{3 \times x}{3 \times 2} - \frac{2 \times (x-2)}{2 \times 3} = .4\]

OpenStudy (sandra):

ok cool np